Solve the system by substitution.
3x - y = -7
x + y = -9

Answer by praseenakos@yahoo.com(507)

QUESTION:

Solve the system by substitution.
3x - y = -7
x + y = -9

ANSWER:

3x - y = -7 -----------(1)

x + y = -9 -----------(2)


Consider equation (2)

x + y = -9

Subtract x from both sides of the above expression.


==> x + y -x = -9 -x


==> x - x + y = -9 - x


==> y = -9 - x


Now substitute this value of y in equation(1)


==> 3x - (-9 - x) = -7

Remove the parenthesis

==> 3x +9 + x = -7


==> 3x + x + 9 = -7


==> 4x + 9 = -7


Subtract 9 from both sides of the above equation.


==> 4x + 9 - 9 = -7 - 9


==> 4x = -16


Divide both sides of the expression by 4

==> 4x/4 = -16/4

==> x = -4

Now substitute this value of x in (2)

Then we have -4 + y = -9

Add 4 on both sides of the equation.


==> -4 + y + 4 = -9 + 4


==> y = -5



To check your answer, plugg these values in the first equation.
Then you can see that these values satisfy equation(1).


So we can say that the solution for the given pair of equations is....

x = -4 and y = -5


Hope you understood.


Regards.

Praseena.